Grade upto college level Physical Chemistry

20% of surface sites are occupied by N2 molecules. The density of surface site is 6.023 * 1014 cm-2 and total surface area is 1000 cm2. The catalyst is heated to 300 K while N2 is completely desorbed into a pressure of 0.001 atm and volume of 2.46 cm3. Find the number of active sites occupied by each N2 molecule.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Hello Student,
Please find the answer to your question
P_{N_{2}}= 0.001 atm, T = 300 K, V = 2.46 cm2
∴ Number of N2 molecules
= PV/RT * NAV = 0.001 *2.46 *10-3/0.0821 *300 * 6.023 * 1023
= 6.016 * 1016
Now total number of surface sites
= Density * Total surface area
= 6.023 * 1014 * 1000 = 6.023 * 1017
Sites occupied by N2 molecules = 20/100 * 6.023 * 1017 = 12.04 * 1016
∴ No. of sites occupied by each N2 molecule
= 12.04 *1016/6.016 * 1016 = 2

Thanks
Deepak patra
askIITians Faculty